OpenStudy (kateqfiofwaejfi):

Show that f(x)=x^2 is continuous for all real numbers (Hint: Using limits --> Lim x-> c- f(x) = lim x--> c+ f(x) = f(c)

OpenStudy (kateqfiofwaejfi):

Answer: Nvm figured it out! limx--> c+ f(x) = c^2 lim x--> c-f(x) = c^2 Therefore lim x--> c f(x) Exists and is c^2 f(c) = c^2 Therefore lim x-->c f(x) = f(c) = c^2

OpenStudy (kateqfiofwaejfi):

So it is continuous for all real numbers